Write a quadratic equation in the variable x having the given numbers as solutions. Type the equation in standard form, ax^2 + bx + c = 0
Problem:
Solution: 2, (only solution)
Thank you!!

Answer by ankor@dixie-net.com(22740)

Write a quadratic equation in the variable x having the given numbers as solutions. Type the equation in standard form, ax^2 + bx + c = 0
Problem:
Solution: 2, (only solution)
:
A solution of 2 will come from factors: (x-2), (x-2)
FOIL (x-2)(x-2) = x^2 - 4x + 4 = 0